OpenStudy (anonymous):

true or false? -6(squared) = (-6)(squared

OpenStudy (saifoo.khan):

Totally F.

OpenStudy (saifoo.khan):

-6^2 = -36 (-6)^2 = 36

OpenStudy (anonymous):

why?

OpenStudy (unklerhaukus):

no \[-(6)^2=-36\] \[(-6)^2=36\]

OpenStudy (anonymous):

oh, i get it

OpenStudy (unklerhaukus):

negative times negative is positive

OpenStudy (saifoo.khan):

-6^2 = -(6*6) = = - 36 (-6)^2 = (-6* -6) = 36